OpenStudy (anonymous):

Joaquin wrote this equation for the area of a parallelogram. 23.7h = 417.12. What is the height (h) of this parallelogram? A. 17.6 units B. 18.1 units C. 393.4 units D. 9885.7 units

OpenStudy (elonasushchik):

The answer is A. cuz all you have to do is divide 23.7 on both sides of the equation 23.7h/23.7 = 417.12/23.7 h = 417.12/23.7

OpenStudy (elonasushchik):

Medal And Fan?! Tx